OpenStudy (anonymous):

f(x) = sqrt(5 - x) g(x) = sqrt(x^2 - 9) (a) f + g = b) f-g=

OpenStudy (anonymous):

\[f[x]+g[x]=\sqrt{5-x}+\sqrt{-9+x^2} \] \[f[x]-g[x]=\sqrt{5-x}-\sqrt{-9+x^2} \]